# Random Pareto front surfaces

### v1

@@ 5.1.2 One-dimensional projection
Parallel coordinates
To visualise the one-dimensional projections of the Pareto front surface, we propose using a parallel coordinates plot, which is a common visualisation tool for high-dimensional Pareto fronts (Tušar and Filipič, 2015). The novelty of our work is that we can also view uncertainty information along each projection as well. For example, in our application we computed and visualised the empirical quantiles and the sample mean obtained using a finite set of samples $\{A^{*}(\boldsymbol{\omega}_{n})\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}\}_{n=1,\dots,N}$. Notably, this overall application is very lightweight to run because all of these statistics can computed and updated very quickly and cheaply on the fly.
5.1.3 Two-dimensional projection
Proposition 5.1 gives us a way to visualise a high-dimensional Pareto front surface by constructing a collection of low-dimensional slices that can be easily visualised instead. To illustrate how one might do this in practice, we created a simple dashboard application that can be used to navigate the set of two-dimensional slices of a Pareto front surface. We give a snapshot of this application in Figure 11. Below we describe the main concept of this application.
Indices and weights
To select the two indices $I=\{i_{1},i_{2}\}$, we used two drop-down lists. The corresponding fixed vector $\mathbf{v}\in\mathcal{V}_{+}^{M-2}$ is then determined by some weight sliders.
Normalisation
Similar to the one-dimensional application, we use an estimate of the objective ranges in order to normalise the objective values and set the reference vector.
Line plot
To visualise the two-dimensional projection of the Pareto front surface, we considered using a regular two-dimensional line plot. In our application, we included visuals on the empirical quantiles and the sample mean of the projected two-dimensional Pareto front surface. Naturally, other useful information could be included as well. For example, to address Remark 5.1, it might be beneficial to also include a parallel coordinates plot of the fixed vector as well—similar to the one used in Figure 10, but only for the fixed components.
Figure 11: An illustration of the dashboard application that we made in order to navigate the space of two-dimensional projections of a Pareto front surface (Section 5.1.3).
5.1.4 Higher dimensional projections
In the previous sections, we have shown how it is possible to view a random Pareto front surface using one-dimensional or two-dimensional slices. The one-dimensional slices gave us information about the marginal performance for each objective. Whilst the two-dimensional slices gave us information about the correlation between any two objectives. Similarly, in order to learn more about the $P$-th order interactions, we need to be able to visualise the $P$-dimensional slices of the Pareto front surface. For example, one might consider adapting one of the methods surveyed by Tušar and Filipič (2015) in order to accomplish this.
@@ 5.2 Uncertainty quantification
We can then appeal to our earlier work in Section 4 in order to study this resulting Pareto front surface distribution. Precisely, we can associate the stochastic process $f:\mathbb{X}\times\Omega\rightarrow\mathbb{R}^{M}$, described in Section 4, with the sampling distribution induced by the latest posterior model: $f(\cdot,\boldsymbol{\omega})\sim p(g|\mathcal{D}_{T})$, where $\boldsymbol{\omega}\in\Omega$ is distributed according to $\mathbb{P}$. We can then analyse and evaluate the posterior Pareto front surface distribution $p(Y_{\boldsymbol{\eta},g}^{*}|\mathcal{D}_{T})$ by studying the corresponding polar parametrised stochastic process described in (10). To showcase how this overall routine works in practice, we present an illustrative example in Section 5.2.1, where we visualise the evolution of a Pareto front surface distribution during a run of Bayesian optimisation. Afterwards, in Section 5.2.2, we then demonstrate how this distributional information can be used in conjunction with the visualisation techniques described in Section 5.1 in order to help us make final decisions.
@@ 5.2.1 Bayesian optimisation
Bayesian optimisation is a popular strategy for black-box optimisation—see the book by Garnett (2023) for a recent overview on this topic. Notably, this experimental design strategy takes advantage of a probabilistic surrogate model in order to determine the best inputs to evaluate. As described above, we can easily take advantage of this probabilistic model in order to compute any Pareto front surface statistic of interest. Practically speaking, we envision that these statistics might be valuable for an active decision maker who is interested in adapting the Bayesian optimisation run for their own purposes. For instance, one might use these statistics in conjunction with the visualisation ideas described in Section 5.1 in order to visualise and better understand the evolution of the Pareto front surface distribution. Given this newfound understanding, a keen decision maker might then actively refine and reprogram [^7] For instance, in the expected hypervolume acquisition criterion (Emmerich et al., 2006), one might use this visual information in order to update the reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. the acquisition procedure in order to target a specific region of interest. Alternatively, earlier work by Binois et al. (2015a) suggested that it might also be possible to use some Pareto front surface statistics, such as the Vorob’ev deviation (22), as a basis for a stopping criteria.
@@
In Figure 12, we illustrate an example of how the Pareto front surface distribution evolves over time during one run of the Bayesian optimisation algorithm applied on the Gaussian mixture model (GMM) (Daulton et al., 2022) and the DTLZ2 (Deb et al., 2002) benchmark problem. For the probabilistic model, we adopted a standard Gaussian process prior (Rasmussen and Williams, 2006) on our objective function and an independent Gaussian observation likelihood on the function observations. For the acquisition function, we used the expected hypervolume improvement (Emmerich et al., 2006). For illustrative convenience, we discretised the input space in both of these benchmark problems to have $|\mathbb{X}|=2^{12}$ points. This latter simplification is only to ensure that we can compute the Pareto front surface of the objective function and the model samples exactly using our polar parametrisation.
@@
There are a number of key observations that we see from the plots in Figure 12. Firstly, we see that the Pareto front surface distribution does indeed slowly converge to the actual Pareto front surface when we observe more data points. This is clearly a desirable property and is something that is expected on these benchmark problems. Secondly, we see that the model Pareto front surface always dominates the sample Pareto front surface in these examples. This makes intuitive sense because the sample front considers only a finite number of $|Y_{T}|=T$ points. In contrast, the model Pareto front surface considers the objective values over the entire input space. Note however that this intuition only holds in our example because the observations were not contaminated with any output noise. Otherwise, when there is observation noise in our data, the sample Pareto front could easily dominate both the actual and model Pareto front surface. For this reason, it is common for practitioners to rely on model-based estimates of the Pareto front surface when the data is noisy. Thirdly, we see that the uncertainty of the Pareto front surface does not necessarily decrease in a monotonic fashion as we add more points. This feature naturally arises in our example because we did not fix the model hyperparameters in our Gaussian process prior. Instead, we followed standard practice and updated these hyperparameters in an online manner by always maximising the latest log marginal likelihood. As a result of this updating step, the overall uncertainty in the Pareto front surface occasionally increased when we incorporate more points—for instance we see this feature when we transition from $T=20$ to $T=30$ in the GMM problem.
@@
Figure 12: An illustration of the change in the Pareto front surface distribution $p(Y_{\boldsymbol{\eta},g}^{*}|\mathcal{D}_{T})$ when we perform Bayesian optimisation on the normalised GMM and DTLZ2 benchmarks. In the plots, we draw the estimated mean front and highlight the region between the $5\%$ and $95\%$ estimated quantile fronts.
@@ 5.2.2 Input decision
where $\boldsymbol{\omega}_{n}\in\Omega$ denotes some independent samples of the random parameter and $\mathbf{y}^{*}\in\mathbb{R}^{M}$ denotes the target vector of interest. Clearly there are many potential loss functions that we can choose in practice. Motivated the work so far, we propose using a length-based loss function because it is interpretable and leads to the desired result.
Length-based loss function
A natural candidate for the loss function would be a frontier loss function (8):
@@
Geometrically, this loss function scores any vector $\mathbf{y}\in\mathbb{R}^{M}$ by computing its projected length along the reference line $L_{\boldsymbol{\eta},\boldsymbol{\lambda}_{\boldsymbol{\eta}}^{*}(\mathbf{y}^{*})}=\{\boldsymbol{\eta}+t\boldsymbol{\lambda}_{\boldsymbol{\eta}}^{*}(\mathbf{y}^{*}):t\in\mathbb{R}\}$ and then comparing it with the desired projected length. We illustrate the efficacy of this strategy for a two-dimensional example in Figure 13. We see clearly in this example that the random vectors $f(\mathbf{x}^{*},\boldsymbol{\omega})\in\mathbb{R}^{M}$ associated with the best inputs $\mathbf{x}^{*}\in\mathbb{X}$ are indeed distributed close to the corresponding target vectors $\mathbf{y}^{*}\in\mathbb{R}^{M}$.
@@ 5.3 Extreme value theory
Extreme value theory is a well-established branch of statistics which studies the distribution of extreme values—see for instance Embrechts et al. (1997); Coles (2001) or Beirlant et al. (2004) for a background on this topic. In this section, we showcase how our polar parametrisation result can be used in order to generalise many existing ideas in this topic to the multivariate setting, where the maximum is defined using the Pareto partial ordering. To the best of our knowledge, the majority of existing work in multivariate extreme value theory has largely focussed on the marginal maximisation setting (Barnett, 1976), where the maximum operation on a set of vectors is defined component-wise. Much less work, if any, has focussed on the setting where the maximum is defined using the Pareto partial ordering. The clear benefit of this latter approach is that it is much more flexible in practice because it can accommodate for the scenario where the various component-wise maximums do not all take place simultaneously.
The most notable results from extreme value theory are the Fisher–Tippett–Gnedenko theorem (Fisher and Tippett, 1928; Gnedenko, 1943) and the Balkema–de Haan–Pickands theorem (Balkema and de Haan, 1974; Pickands, 1975). The former result is concerned with the asymptotic distribution of the maximum order statistic associated with a collection of independent and identically distributed univariate random variables. In words, it states that the distribution of the maximum, upon proper normalisation, can only converge in distribution to either a Gumbel, Fréchet or Weibull distribution. In contrast, the latter result is concerned with the limiting distribution of the corresponding conditional excess distribution (Definition 5.2). Conceptually, this result tells us that the distribution of the tail events, pass some threshold, can be closely approximated by a generalised Pareto distribution. For reference purposes, we repeat both of these results below in Theorem 5.1 and Theorem 5.2, respectively.
Definition 5.1 (Maximum domain of attraction)
Theorem 5.1 (Extreme value distribution)
Definition 5.2 (Conditional excess distribution)
$F_{u}(x):=\mathbb{P}[Y-u\leq x|Y>u]=\frac{F(u+x)-F(u)}{1-F(u)},$
Theorem 5.2 (Generalised Pareto distribution)
@@
The existing work on multivariate extreme value theory has largely focussed on the marginal maximisation setting where the maximum of a collection of $N$ independent and identically-distributed vectors $Y_{1},\dots,Y_{N}\in\mathbb{R}^{M}$ are defined component-wise. The traditional goal of interest is then to study the multivariate generalisation of the MDA for the resulting multivariate distribution function $F:\mathbb{R}^{M}\rightarrow\mathbb{R}$,
@@
for $\mathbf{x}\in\mathbb{R}^{M}$—see Beirlant et al. (2004, Chapter 8) for an overview on this topic. The primary benefit of using this component-wise definition is that we can immediately apply both Theorem 5.1 and Theorem 5.2 in order to determine the asymptotic properties of the corresponding marginal distributions. The remaining challenge is then to identify the corresponding dependence structure between the components. This latter problem turns out to be a major hurdle in multivariate extreme value theory because this dependency structure cannot necessarily be described using a finitely parametrised model. Notably, the estimation and analysis of this dependency structure is still an area of active research—see Beirlant et al. (2004, Chapter 8) for an in-depth discussion.
@@
To the best of our knowledge, there does not exist much work regarding the study of multivariate extreme value theory when we use the Pareto definition of the vector-valued maximum (Definition 2.4). We attribute this lack of interest based on the fact that the Pareto maximum is very challenging to work with in practice. One remarkable outcome of our work is that it manages to connect the definition of the Pareto maximum with the definition of the component-wise maximum. Specifically, our polar parametrisation result tells us that the Pareto maximum can be completely characterised by its projected length process (11), which is defined as an infinite collection of scalarised maximums. To put it more concretely, if we had a collection of identically-distributed vectors $Y_{1},\dots,Y_{N}\in\mathbb{R}^{M}$, then the corresponding Pareto maximum is governed by the following projected length process
$\mathcal{L}_{\boldsymbol{\eta}}=\{\max(\{s_{\boldsymbol{\eta},\boldsymbol{\lambda}}(Y_{1}),\dots,s_{\boldsymbol{\eta},\boldsymbol{\lambda}}(Y_{N})\})\geq 0:\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}\}$
for any reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. As a consequence of this observation, many of the results that are known for the component-wise maximum can now be adapted to the Pareto maximum setting. For example, we can use Theorem 5.1 to determine the possible asymptotic distributions of the projected length process along each positive direction. Similarly, we can use Theorem 5.2 in order to approximate the distribution of the tails of the projected length process conditional on the fact that it dominates some specified polar surface. That is, instead of setting just a single threshold value of $u\in\mathbb{R}$, we set the threshold to be a polar surface defined by some set $\mathcal{U}=\{u_{\boldsymbol{\lambda}}\geq 0:\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}\}$.
As a proof of concept for these general ideas, we present a simple example in Proposition 5.2, where we study the asymptotic distribution of a Pareto front surface constructed using a collection of independent Weibull distributed vectors. Formally, this result states that, upon proper normalisation, the projected length process of this Pareto front surface distribution converges in distribution to a Gumbel distribution along each positive direction—the proof of this result is presented in Section A.11. To illustrate this convergence, we present some visual two-dimensional examples in Figure 14. In these plots, we fixed [^9] The rate parameter only controls the relative scaling of each objective and therefore it does not play a major role in the assessment of the empirical result. the rate parameter $\beta^{(m)}>0$ and varied the choice of shape parameter $\alpha>0$. On the whole, we see that the empirical and limiting distributions are already quite similar even at small sample sizes such as $N=64$. Conceptually, we see that varying the shape parameter $\alpha>0$ amounts to changing from a concave Pareto front surface distribution when $\alpha\in(0,1)$ to a convex Pareto front surface distribution when $\alpha>1$.
Proposition 5.2 can also be used in conjunction with Theorem 5.2. In words, this result tells us that the tails of the corresponding projected length process are approximately distributed according to an exponential distribution. To demonstrate this property, we present an illustrative example in Figure 15, where we study the conditional excess probabilities associated with the $M=2$ dimensional Pareto front surface distributions described in Figure 14. That is, we plot the probabilities
@@
Remark 5.3 (Endpoint estimation)
As identified by Binois et al. (2015b), the Pareto front surface of an objective function $g:\mathbb{X}\rightarrow\mathbb{R}^{M}$ can also be viewed as the uppermost quantile of a Pareto front surface distribution (Binois et al., 2015b, Theorem 2.1),
$\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\{g(\mathbf{x})\}_{\mathbf{x}\in\mathbb{X}}]=\lim_{\alpha\rightarrow 1}\mathcal{Q}_{\mathbf{x}\sim p(\mathbf{x})}[\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\{g(\mathbf{x})\}],\alpha],$
where $p(\mathbf{x})\geq 0$ denotes the density of any distribution over the input space whose essential support covers the entire input space; for instance, a uniform distribution over $\mathbb{X}$. As emphasised by Binois et al. (2015b), the clear benefit of this result is that it transforms the vector-valued optimisation problem into an endpoint estimation problem. This latter topic is naturally related with extreme value theory (Hall, 1982; Loh, 1984; Girard et al., 2012).

### v2

@@ 5.1.2 Dashboard applications
To visualise the one-dimensional projections of the Pareto front surface, we propose using a parallel coordinates plot, which is a common visualisation tool for high-dimensional Pareto fronts (Tušar and Filipič, 2015). The novelty of our work is that we can also view uncertainty information along each projection as well. For example, in our application we computed and visualised the empirical quantiles and the sample mean obtained using a finite set of samples $\{A^{*}(\boldsymbol{\omega}_{n})\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}\}_{n=1}^{N}$. Notably, this overall application is very lightweight to run because all of these statistics can computed and updated very quickly and cheaply on the fly.
Figure 10: An illustration of the dashboard application that we made in order to navigate the space of one-dimensional projections of a Pareto front surface.
Two-dimensional projection
Our two-dimensional dashboard application works in very much the same way as our one-dimensional application described above. The only key difference now is the introduction of two additional drop-down lists, which are used to determine the indices of the set $I=\{i_{1},i_{2}\}$. The corresponding fixed vector $\mathbf{v}\in\mathcal{V}_{+}^{M-2}$ is then determined by the remaining (normalised) weight sliders.
To visualise the two-dimensional slices of the Pareto front surface distribution, we considered using a regular two-dimensional line plot. In our application, we included visuals on the empirical quantiles and the sample mean of the projected two-dimensional Pareto front surface. Naturally, other useful information could be included as well. For example, to address Remark 5.1, it might be beneficial to also include a parallel coordinates plot of the fixed vector as well—similar to the one used in Figure 10, but only for the fixed components.
Figure 11: An illustration of the dashboard application that we made in order to navigate the space of two-dimensional projections of a Pareto front surface.
Higher dimensional projections
The one-dimensional slices gave us information about the marginal performance for each objective. Whilst the two-dimensional slices gave us information about the correlation between any two objectives. Similarly, in order to learn more about the $P$-th order interactions, we need to be able to visualise the $P$-dimensional slices of the Pareto front surface. For example, one might consider adapting one of the methods surveyed by Tušar and Filipič (2015) in order to accomplish this.
@@ 5.2 Uncertainty quantification
We can then appeal to our earlier work in Section 4 in order to study this resulting Pareto front surface distribution. More precisely, we can associate the random function $f:\mathbb{X}\times\Omega\rightarrow\mathbb{R}^{M}$, described in Section 4, with the sampling distribution induced by the latest posterior model: $f(\cdot,\boldsymbol{\omega})\sim p(g(\cdot)|\mathcal{D}_{T})$, where $\boldsymbol{\omega}\in\Omega$ is distributed according to $\mathbb{P}$. We can then analyse and evaluate the posterior Pareto front surface distribution $p(Y_{\boldsymbol{\eta},g}^{*}|\mathcal{D}_{T})$ by studying the corresponding polar parameterised random variable described in (10). To showcase how this overall routine works in practice, we present an illustrative example in Section 5.2.1, where we visualise the evolution of a Pareto front surface distribution during different runs of Bayesian optimisation. Afterwards, in Section 5.2.2, we then demonstrate how this distributional information can be used in conjunction with the visualisation techniques described in Section 5.1 in order to help us make final decisions.
@@ 5.2.1 Bayesian optimisation
Bayesian optimisation is a popular strategy for black-box optimisation—see the book by Garnett (2023) for a recent overview on this topic. Notably, this experimental design strategy takes advantage of a probabilistic surrogate model in order to determine the best inputs to evaluate. As described above, we can easily take advantage of this probabilistic model in order to compute any Pareto front surface statistic of interest. Practically speaking, we envision that these statistics might be valuable for an active decision maker who is interested in adapting the Bayesian optimisation run for their own purposes. For instance, one might use these statistics in conjunction with the visualisation ideas described in Section 5.1 in order to visualise and better understand the evolution of the Pareto front surface distribution. Given this newfound understanding, a keen decision maker might then actively refine and reprogram [^7] For instance, in the expected hypervolume acquisition criterion (Emmerich et al., 2006), one might use this visual information in order to update the reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. the acquisition procedure in order to target a specific region of interest. Alternatively, earlier work by Binois et al. (2015) suggested that it might also be possible to use some Pareto front surface statistics, such as the Vorob’ev deviation (22), as a basis for a stopping criteria.
@@
In Figure 12, we illustrate an example of how the Pareto front surface distribution evolves over time during one run of the Bayesian optimisation algorithm applied on the Gaussian mixture model (GMM) (Daulton et al., 2022) and the DTLZ2 (Deb et al., 2002) benchmark problem. For the probabilistic model, we adopted a standard Gaussian process prior (Rasmussen and Williams, 2006) on our objective function and an independent Gaussian observation likelihood on the function observations. For the acquisition function, we used the expected hypervolume improvement (Emmerich et al., 2006). For illustrative convenience, we discretised the input space in both of these benchmark problems to have $|\mathbb{X}|=2^{12}$ points. This latter simplification is only to ensure that we can compute the Pareto front surface of the objective function and the model samples exactly using our polar parameterisation.
@@
There are a number of key observations that we see from the plots in Figure 12. Firstly, we see that the Pareto front surface distribution does indeed slowly converge to the actual Pareto front surface when we observe more data points. This is clearly a desirable property and is something that is expected on these benchmark problems. Secondly, we see that the model Pareto front surface always dominates the sample Pareto front surface in these examples. This makes intuitive sense because the sample front considers only a finite number of $|Y_{T}|=T$ points. In contrast, the model Pareto front surface considers the objective values over the entire input space. Note however that this intuition only holds in our example because the observations were not contaminated with any output noise. Otherwise, when there is observation noise in our data, the sample Pareto front could easily dominate both the actual and model Pareto front surface. For this reason, it is common for practitioners to rely on model-based estimates of the Pareto front surface when the data is noisy. Thirdly, we see that the uncertainty of the Pareto front surface does not necessarily decrease in a monotonic fashion as we add more points. This feature naturally arises in our example because we did not fix the model hyperparameters in our Gaussian process prior. Instead, we followed standard practice and updated these hyperparameters in an online manner by always maximising the latest log marginal likelihood (Balandat et al., 2020). As a result of this updating step, the overall uncertainty in the Pareto front surface occasionally increased when we incorporate more points—for instance we see this feature when we transition from $T=20$ to $T=30$ in the GMM problem.
@@
Figure 12: An illustration of the change in the Pareto front surface distribution $p(Y_{\boldsymbol{\eta},g}^{*}|\mathcal{D}_{T})$ when we perform Bayesian optimisation on the normalised GMM and DTLZ2 benchmarks. In the plots, we draw the estimated mean front and highlight the region between the $5\%$ and $95\%$ estimated quantile fronts in blue.
@@ 5.2.2 Input decision
where $\{\boldsymbol{\omega}_{n}\}_{n=1}^{N}\in\Omega$ denotes i.i.d. samples of the random parameter and $\mathbf{y}^{*}\in\mathbb{R}^{M}$ denotes the target vector of interest. Clearly there are many potential loss functions that we can choose in practice. Motivated by the work so far, a natural candidate for the loss function would be a frontier loss function (8):
@@
Geometrically, this loss function scores any vector $\mathbf{y}\in\mathbb{R}^{M}$ by computing its projected length along the reference line $L_{\boldsymbol{\eta},\boldsymbol{\lambda}_{\boldsymbol{\eta}}^{*}(\mathbf{y}^{*})}=\{\boldsymbol{\eta}+t\boldsymbol{\lambda}_{\boldsymbol{\eta}}^{*}(\mathbf{y}^{*}):t\in\mathbb{R}\}$ and then comparing it with the desired projected length. We illustrate the efficacy of this strategy for a two-dimensional example in Figure 13 based on the four bar truss optimisation problem (Cheng and Li, 1999; Tanabe and Ishibuchi, 2020). We see clearly in this example that the random vectors $f(\mathbf{x}^{*},\boldsymbol{\omega})\in\mathbb{R}^{M}$ associated with the best inputs $\mathbf{x}^{*}\in\mathbb{X}$ are indeed distributed close to the corresponding target vectors $\mathbf{y}^{*}\in\mathbb{R}^{M}$.
@@ 5.3 Extreme value theory
Extreme value theory is a well-established branch of statistics which studies the distribution of extreme values—see for instance Embrechts et al. (1997); Coles (2001) or Beirlant et al. (2004) for a background on this topic. In this section, we showcase how our polar parameterisation result can be used in order to generalise many existing ideas in this topic to the multivariate setting, where the maximum is defined using the Pareto partial ordering. To the best of our knowledge, the majority of existing work in multivariate extreme value theory has largely focussed on the marginal maximisation setting (Barnett, 1976), where the maximum operation on a set of vectors is defined component-wise. No work has focussed on the setting where the maximum is defined using the Pareto partial ordering. We attribute this lack of interest based on the fact that the Pareto maximum is very challenging to work with in practice. Despite this, there are clear benefits for adopting the latter approach over the former. Most notably, the Pareto approach is much more flexible because it can also accommodate for the scenario where the various component-wise maximums do not all take place simultaneously.
The most notable results from extreme value theory are the Fisher–Tippett–Gnedenko theorem (Fisher and Tippett, 1928; Gnedenko, 1943) and the Balkema–de Haan–Pickands theorem (Balkema and de Haan, 1974; Pickands, 1975). The former result is concerned with the asymptotic distribution of the maximum order statistic associated with a collection of independent and identically distributed univariate random variables. In words, it states that the distribution of the maximum, upon proper normalisation, can only converge in distribution to either a Gumbel, Fréchet or Weibull distribution. In contrast, the latter result is concerned with the limiting distribution of the corresponding conditional excess distribution (Definition A.2). Conceptually, this result tells us that the distribution of the tail events, pass some threshold, can be closely approximated by a generalised Pareto distribution. For reference purposes, we recall both of these results in the Appendix—namely, Theorem A.1 and Theorem A.2, respectively.
@@
The existing work on multivariate extreme value theory has largely focussed on the marginal maximisation setting where the maximum of a collection of $N$ independent and identically-distributed vectors $Y_{1},\dots,Y_{N}\in\mathbb{R}^{M}$ are defined component-wise. The traditional goal of interest is then to study the multivariate generalisation of the maximum domain of attraction (MDA) for the resulting multivariate distribution function $F:\mathbb{R}^{M}\rightarrow\mathbb{R}$,
@@
for $\mathbf{x}\in\mathbb{R}^{M}$—see Beirlant et al. (2004, Chapter 8) for an overview on this topic. The primary benefit of using this component-wise definition is that we can immediately apply both Theorem A.1 and Theorem A.2 in order to determine the asymptotic properties of the corresponding marginal distributions. The remaining challenge is then to identify the corresponding dependence structure between the components. This latter problem turns out to be a major hurdle in multivariate extreme value theory because this dependency structure cannot necessarily be described using a finitely parameterised model. Notably, the estimation and analysis of this dependency structure is still an area of active research—see Beirlant et al. (2004, Chapter 8) for an in-depth discussion.
@@
One notable outcome of our work is that it manages to connect the definition of the Pareto maximum with the definition of the component-wise maximum. Specifically, our polar parameterisation result tells us that the Pareto maximum can be completely characterised by its projected lengths (11), which is defined as an infinite collection of scalarised maximums. To put it more concretely, if we had a collection of identically-distributed vectors $Y_{1},\dots,Y_{N}\in\mathbb{R}^{M}$, then the corresponding Pareto maximum is governed by the following infinite-dimensional random variable
$\mathcal{L}_{\boldsymbol{\eta}}=\{\max(\{s_{\boldsymbol{\eta},\boldsymbol{\lambda}}(Y_{1}),\dots,s_{\boldsymbol{\eta},\boldsymbol{\lambda}}(Y_{N})\})\geq 0:\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}\},$
for any reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. As a consequence of this observation, many of the results that are known for the component-wise maximum can now be adapted to the Pareto maximum setting. For example, we can use Theorem A.1 to determine the possible asymptotic distributions of the projected length process along each positive direction. Similarly, we can use Theorem A.2 in order to approximate the distribution of the tails of the projected length process conditional on the fact that it dominates some specified polar surface. That is, instead of setting just a single threshold value of $u\in\mathbb{R}$, we set the threshold to be a polar surface defined by some set $\mathcal{U}=\{u_{\boldsymbol{\lambda}}\geq 0:\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}\}$.
As a proof of concept for these general ideas, we present a simple example in the proposition below, where we study the asymptotic distribution of a Pareto front surface constructed using a collection of independent Weibull distributed vectors—the proof of this result is presented in Section A.9.1.
@@
To illustrate this convergence, we present some visual two-dimensional examples in Figure 14. In these plots, we fixed [^8] The rate parameter only controls the relative scaling of each objective and therefore it does not play a major role in the assessment of the empirical result. the rate parameter $\beta^{(m)}>0$ and varied the choice of shape parameter $\alpha>0$. On the whole, we see that the empirical and limiting distributions are already quite similar even at small sample sizes such as $N=64$. Conceptually, we see that varying the shape parameter $\alpha>0$ amounts to changing from a concave Pareto front surface distribution when $\alpha\in(0,1)$ to a convex Pareto front surface distribution when $\alpha>1$.
Proposition 5.2 can also be used in conjunction with Theorem A.2. In words, this result tells us that the tails of the corresponding projected length process are approximately distributed according to an exponential distribution. To demonstrate this property, we present an illustrative example in Figure 15, where we study the conditional excess probabilities associated with the $M=2$ dimensional Pareto front surface distributions described in Figure 14. That is, we plot the probabilities